OpenStudy (anonymous):

5+5=?????????? im confused need help last question of tha day :)

OpenStudy (anonymous):

Ok so if i have 5 dollars , right? and you give me 5 more dollars? what would that be?

OpenStudy (anonymous):

pleas mods no warning this is real

OpenStudy (anonymous):

Look ready. 1+1+1+1+1+1+1+1+1+1 count that please

OpenStudy (anonymous):

Look at your fingers.

OpenStudy (anonymous):

ok i m looking

OpenStudy (lgbasallote):

\[\huge 5 + 5 = \int \limits_0^2 (5x)\text dx\]

OpenStudy (anonymous):

how many do you have on your left hand?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

how many do you have on your right hand?

OpenStudy (anonymous):

Windows: Start>All Programs>Accessories>Calculator Macintosh: Command+Space>Type "calculator">Hit enter. Also: Google 5+5 This works for ANY arithmetic expression!!

OpenStudy (anonymous):

4 one was amputateed

OpenStudy (anonymous):

right. so what's the objective of your assignment?

OpenStudy (anonymous):

Oh man im sorry to here that :( . You probably made him feeel bad because he doesnt havve 10 fingers like use :(

OpenStudy (lgbasallote):

you can also use the formula \[\huge \int \limits_0^{\frac{\pi}{2}} \mathcal L \left \{ \frac{d}{dx}\left(\frac{\Gamma \delta}{\Gamma \alpha}\right) \right\}\]

OpenStudy (anonymous):

trying to figure out how much is 5+5 oh dam i missd numbers its 5+5/by the sqr of 3 and then crooss multyply and at the end subtract 10

OpenStudy (anonymous):

I'm so sure. So Alex, you can go ahead and look at your hand again, curl it up like fist and knock it on your forehead 10 times.

OpenStudy (lgbasallote):

use the formula i just gave

OpenStudy (anonymous):

woah woah excuse me thats inappropriate EXCUSE ME MOD CAN YOU REPORT THIS WOMEN

OpenStudy (anonymous):

ok and dont be mean miss u hurt my feelings

OpenStudy (anonymous):

ok formula

OpenStudy (anonymous):

oh wait almost got it

OpenStudy (anonymous):

In my day, trolling actually meant something.

OpenStudy (anonymous):

Wait man look at you question how many question marks do you have xD

OpenStudy (anonymous):

WHATS TROLLING ok i got the answer

OpenStudy (anonymous):

its cero thanks every one and i have 10 on my feet

OpenStudy (anonymous):

Good Job :)

OpenStudy (anonymous):

so have a daarling night people may the force be with u

OpenStudy (anonymous):

No.. may the force be with you alex.. my young padiwan

OpenStudy (lgbasallote):

@Herp_Derp that wasnt a real formula...those were just random symbols i learned the past year that i put in one thingy

OpenStudy (anonymous):

lol